Clint is constructing two adjacent rectangular dog pens. Each pen will be three times as long as it is wide, and the pens will share a common long side. If clint has 65 ft of fencing, what are the dimensions of each pen?
Found 3 solutions by mananth, ikleyn, greenestamps:
Answer by mananth(16949)

Clint is constructing two adjacent rectangular dog pens. Each pen will be three times as long as it is wide, and the pens will share a common long side. If clint has 65 ft of fencing, what are the dimensions of each pen?
Let width be x
Length =3x
Perimeter =3x + 3x + 3x + 3x + x + x = 14x
14x=65
X=4.64 ft
Length= 13.92ft
mananth@hotmail.com

Answer by ikleyn(53996)

.
Clint is constructing two adjacent rectangular dog pens. Each pen will be three times as long as it is wide,
and the pens will share a common long side. If clint has 65 ft of fencing, what are the dimensions of each pen?
~~~~~~~~~~~~~~~~~~~~~~~~~


        The solution in the post by @mananth is incorrect conceptually: he incorrectly setup his governing equation.
        I came to bring a correct solution.


Let x be width of the pens, in feet.
Then their length is 3x feet, according to the problem.


We have 3 long sides of the length 3x ft each, and 4 short sides of the length x ft each.


So, the total length of all dimensions is   3*(3x) + 4x.

Therefore, the equation for the total fence length is

    3*(3x) + 4x = 65  feet.


Simplify and find x

    13x = 65,

      x = 65/13 = 5 feet.


ANSWER.  The dimensions of each pen are  5 ft x 15 ft.

Solved correctly.

--------------------------


Thanks to tutor @greenestamps for noticing my typo.
I just fixed it.




Answer by greenestamps(13374)



For tutor @ikleyn....

You have a typo in your response that might confuse students. There are 4 "sides" of length x, not 3.